OpenStudy (anonymous):

I WILL GIVE MEDAL AND FAN!!! What is the value of b? b/-3=1.5 b =

OpenStudy (alexandervonhumboldt2):

mutply both sides by -3: you will get b=1.5*(-3)=-(1.5*3)=??

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

No one want medal?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (dm91):

your answer is b=-4.5

OpenStudy (anonymous):

@dm91 are you sure?

OpenStudy (anonymous):

I will give you a fan if you help me with two more. I already gave you a medal

OpenStudy (dm91):

yes because you have to do the opposite of division and that is multiplication so you have to go -3*-3 and you can't do it so then you have to do 1.5*-3=-4.5 and you are trying to find out what B is so 1.5*-3=-4.5 so -4.5 would be your answer

OpenStudy (dm91):

ok i will help you but close this question and open a new one Openstudy rules

OpenStudy (anonymous):

ok

OpenStudy (dm91):

tag me in it and i will come